# OTCliM: generating a near-surface climatology of optical turbulence strength ($C_{n}^{2}$) using gradient boosting

### v1

@@ IV.1 New York State Mesonet (NYSM)
Here, the subscripts p25 and p75 indicate the 25th and 75th percentiles of the original distribution used for scaling.
@@ IV.2 ERA5 reanalysis data
The ERA5 reanalysis [30] serves as the input dataset for OTCliM from which $C_{n}^{2}$ shall be estimated. ERA5 is available globally from 1950 until the present and, thus, includes our NYSM training locations and times. We extract ERA5 input data for all NYSM stations and training years at the ERA5 grid points closest to the respective stations. ERA5 grid points represent 1/4-degree grid boxes (ca. 30 km by 30 km), so by selecting the closest point, we obtain data representative for the grid box containing the station. Figure 2 depicts the location of the flux stations (grey markers) and their collocated grid boxes (orange square) and also gives an impression of how the coarse ERA5 represents (a) land, sea, and (b) terrain.
@@
A complete list of the ERA5 variables selected as features is presented in Tab. 2. Many listed features are (partially) redundant and/or (partially) correlated. Since it is initially unknown which features are most suitable for estimating near-surface $C_{n}^{2}$, we aim for a broad feature space. During training, the GBM algorithm will identify and base its predictions on the most important features. Through the post-training feature importance analysis, we can identify which features the trained model deems relevant and assess their physical relevance. It is usually preferred in ML to minimize the number of features and reduce model complexity. However, for tree-based algorithms like GBM used in this study, irrelevant features are automatically mostly disregarded, mitigating this concern.
The features in Tab 2 without ERA5 variable names are so-called engineered features, meaning that they are derived from one or more ERA5 variables. In the following, we detail the variable selection and the feature engineering.
@@ V Results: feature importance
The three urban stations, BKLN, QUEE, and STAT, and the coastal SOUT site deviate most obviously from the mean FI distribution. In particular, BKLN and QUEE show an above-average dependency of estimated $C_{n}^{2}$ on wind direction. Both stations are located on rooftops and have some tall buildings in their neighborhood [36]. Such inhomogeneous urban conditions are known to influence local measurements [37], so it seems reasonable that turbulence strength observed at BKLN and QUEE depends more strongly on wind direction compared to other more rural locations. All four stations (urban + SOUT) also show below-average dependency on $Q_{H}$, which we view as a feature capturing the diurnal cycle. Instead, the models seem to have picked $\Sigma\,\text{hr}^{\prime}$ or $h_{i}$ as predictors for this information. This behavior differs from all other models, where $Q_{H}$ is typically the second most important input, in line with more traditional physics-based $C_{n}^{2}$ models, such as W71. This shift, however, need not be viewed as unfavorable but as a demonstration of the power of ML-based modeling to shift from traditional to “unconventional” features if complex flow conditions require it. The downside is that we expect such models to perform more poorly when applied to other sites where the more traditional features are relevant. Especially, BKLN is expected to generalize poorly because it also depends strongly on the temperature features $T_{\square}$, deemed irrelevant by almost all other models.
@@ VI Results: OTCliM performance
Figure 5: Cross-site evaluation performance of OTCliM models. The row of a heatmap presents the performance of the model trained on site $s$ when evaluated on all other sites $\tilde{s}\in\mathcal{S}$. The left heatmaps show the performance degradation compared to the MCP case, i.e, the relative scores $\square_{s\rightarrow\tilde{s}}^{\prime}=\square_{s\rightarrow\tilde{s}}/\square_{s\rightarrow{s}}$, and the right ones show the absolute scores for reference. The histograms depict the distributions of both heatmaps.
VI.1 Temporal extrapolation
@@
The insight that the training year selection has a low influence on the model performance has high practical relevance. It suggests that models can be trained on archive data and do not necessarily require recent observations iff the site’s climatology remains comparable between the training period and the envisioned period of model application. Drastic changes in the station’s surroundings, e.g., through construction projects or climate change, would likely break this assumption and, thus, the model’s applicability.
@@ Acknowledgements.
This publication is part of the project FREE – Optical Wireless Superhighways: Free photons (at home and in space) (with project number P19-13) of the research programme TTW-Perspectief, which is (partly) financed by the Dutch Research Council (NWO). This research is made possible by the New York State (NYS) Mesonet. Original funding for the NYS Mesonet (NYSM) buildup was provided by Federal Emergency Management Agency grant FEMA-4085-DR-NY. The continued operation and maintenance of the NYSM is supported by National Mesonet Program, University at Albany, Federal and private grants, and others.
Appendix A Details about baseline models

### v2

@@ IV.1.3 Distribution scaling
Here, the subscripts p25 and p75 indicate the 25th and 75th percentiles of the original distribution used for scaling. More details are given in Appendix B.
@@ IV.2 ERA5 reanalysis data
The ERA5 reanalysis [37] serves as the input dataset for OTCliM from which $C_{n}^{2}$ shall be estimated. ERA5 is available globally from 1950 until the present and, thus, includes our NYSM training locations and times. We extract ERA5 input data for all NYSM stations and training years at the ERA5 grid points closest to the respective stations. ERA5 grid points represent 1/4-degree grid boxes (ca. 30 km by 30 km), so by selecting the closest point, we obtain data representative for the grid box containing the station. Figure 2 depicts the location of the flux stations (grey markers) and their collocated grid boxes (orange square) and also gives an impression of how the coarse ERA5 represents (a) land, sea, and (b) terrain. The temporal resolution/sampling rate also differs between ERA5 (1 h) and the observed $C_{n}^{2}$ dataset (5 min). ERA5 data are instantaneous snapshots available at each full hour, so we match each ERA5 sample with the average of the six 5-min $C_{n}^{2}$ values ±$\pm 15$ min around the hour. This 30 min average centered around the hour smoothes the $C_{n}^{2}$ signal to counteract potential temporal misalignment between observations and reanalysis.
@@
A complete list of the ERA5 variables selected as features is presented in Tab. 2. Many listed features are (partially) redundant and/or (partially) correlated. Since it is initially unknown which features are most suitable for estimating near-surface $C_{n}^{2}$, we aim for a broad feature space. During training, the GBM algorithm will identify and base its predictions on the most important features. Through the post-training feature importance analysis, we can identify which features the trained model deems relevant and assess their physical relevance. It is usually preferred in ML to minimize the number of features and reduce model complexity. However, this concern does not apply to tree-based algorithms like GBM used in this study because such algorithms only consider features during fitting, which increase the model accuracy. Consequently, features contributing not or little to the model’s accuracy are not selected (for splitting) while the trees are grown [39], so their impact on model complexity is low [52].
The features in Tab. 2 without ERA5 variable names are so-called engineered features, meaning that they are derived from one or more ERA5 variables. In the following, we detail the variable selection and the feature engineering.
@@ V Results: feature importance
The three urban stations, BKLN, QUEE, and STAT, and the coastal SOUT site deviate most obviously from the mean FI distribution. In particular, BKLN and QUEE show an above-average dependency of estimated $C_{n}^{2}$ on wind direction. Both stations are located on rooftops and have some tall buildings in their neighborhood [43]. Such inhomogeneous urban conditions are known to influence local measurements [44], so it seems reasonable that turbulence strength observed at BKLN and QUEE depends more strongly on wind direction compared to other more rural locations. All four stations (urban + SOUT) also show below-average dependency on $Q_{H}$, which we view as a feature capturing the diurnal cycle. Instead, the models seem to have picked $\Sigma\,\text{hr}^{\prime}$ or $h_{i}$ as predictors for this information. This behavior differs from all other models, where $Q_{H}$ is typically the second most important input, in line with more traditional physics-based $C_{n}^{2}$ models, such as W71. This shift, however, need not be viewed as unfavorable but as a demonstration of the power of ML-based modeling to shift from traditional to “unconventional” features if complex flow conditions require it. The downside is that we expect such models to perform more poorly when applied to other sites where the more traditional features are relevant. Especially, BKLN is expected to generalize poorly because it also depends strongly on the temperature features $T_{2}$, $T_{d2}$, and $T_{soil}$, deemed irrelevant by almost all other models.
@@ VI.1 Temporal extrapolation
The insight that the training year selection has a low influence on the model performance has high practical relevance. It suggests that models can be trained on archive data and do not necessarily require recent observations if and only if the site’s climatology remains comparable between the training period and the envisioned period of model application. Drastic changes in the station’s surroundings, e.g., through construction projects or climate change, would likely break this assumption and, thus, the model’s applicability.
@@ VI.2 Cross-site evaluation
Figure 5: Cross-site evaluation performance of OTCliM models. The rows of the heatmaps present the performance of the models trained on site $s$ when evaluated on all other sites $\tilde{s}\in\mathcal{S}$. The left heatmaps show the performance degradation compared to the MCP case, i.e, the relative scores $\square_{s\rightarrow\tilde{s}}^{\prime}=\square_{s\rightarrow\tilde{s}}/\square_{s\rightarrow{s}}$ with $\square$ representing $r$ or $\epsilon$, and the right ones show the absolute scores for reference. The histograms depict the distributions of both heatmaps.
@@ Acknowledgements.
This publication is part of the project FREE – Optical Wireless Superhighways: Free photons (at home and in space) (with project number P19-13) of the research programme TTW-Perspectief, which is (partly) financed by the Dutch Research Council (NWO). This research is made possible by the New York State (NYS) Mesonet. Original funding for the NYS Mesonet (NYSM) buildup was provided by Federal Emergency Management Agency grant FEMA-4085-DR-NY. The continued operation and maintenance of the NYSM is supported by National Mesonet Program, University at Albany, Federal and private grants, and others. Sukanta Basu is grateful for financial support from the State University of New York’s Empire Innovation Program.
Appendix A Ratio of stable to unstable conditions in dataset before and after QA/QC
Figure 6: Distribution of bulk potential temperature gradient $\Gamma=(\theta_{9}-\theta_{2})/7\,\text{m}$ before (blue) and after (orange) applying the quality assurance and control (QA/QC) steps presented in §IV.1.2. The split between unstable ($\Gamma<0$, solid bar) and stable ($\Gamma>0$, hatched bar) atmospheric conditions is visualized by the bar charts in each panel or quantified in the respective legend. The QUEE site is omitted fully because of a malfunctioning instrument.
A three-step quality assurance and control (QA/QC) procedure is described in §IV.1.2, which aims at filtering unphysical values from our $C_{n}^{2}$ dataset to lay the foundation for high-quality OTCliM models. As discussed in the respective section, the QA/QC procedure discards approximately half of the data points across all stations and all years. We aim to not significantly change the ratio of unstable to stable atmospheric conditions by QA/QC to expose the OTCliM models to turbulence conditions representative of the respective site. In Fig. 6, we present the distribution of atmospheric stability before and after QA/QC for each location. Stability is quantified through the bulk potential temperature gradient $\Gamma=(\theta_{9}-\theta_{2})/7\,\text{m}$, which is computed from the potential temperature $\theta$ measured by two thermometers at 2 m and 9 m above ground [49]. These instruments differ from the sonic anemometers used to estimate $C_{n}^{2}$. They are more reliable, resulting in only 1.5% of missing values on average due to instrumentation problems compared to 16% for the sonic anemometers. The only exception is the QUEE site, which did not have any 9 m observations during our time of interest, which is why it is omitted in this discussion. Figure 6 shows that the unstable/stable ratio (see bar chart insets) does not change much at most sites due to quality control. In general, the number of stable condition samples decreases because stable conditions occur more frequently in winter when instrument malfunctioning due to rain, ice, or snow is more common. Additionally, the inertial range shrinks with increasing stability and can even disappear [60], resulting in more failed fits of the 2/3 slope (QA/QC step one). Since the balance between stable and unstable cases remains better than approx. 30/70 for all stations, we view our dataset as reasonably well-balanced and adequate for training.
Appendix B Scaling of $\log_{10}C_{n}^{2}$ target data
As described in Sect. IV.1.3 of the main text, the $\log_{10}C_{n}^{2}$ data are normalized using the 25th and 75th percentiles of the site-specific $\log_{10}C_{n}^{2}$ training data (cf. Eq. 10). The effect of this scaling is illustrated in Fig. 7, which presents the unscaled (left) and scaled (right) data colored by location. The plots demonstrate that the scaling makes the range of $\log_{10}C_{n}^{2}$ more comparable between sites, which is crucial to obtain comparable performance and FI scores. That is especially important for the urban BKLN site (dark blue), where the unscaled right tail of the distribution is shifted to the right by half an order of magnitude (higher OT strength) compared to the other sites. To keep the scaling robust and general (i.e., applicable to non-gaussian distributions), we scale using the inter-quartile range and not the commonly used standard deviation.
Figure 7: Comparison between distributions of unscaled (left) and scaled (right) $\log_{10}C_{n}^{2}$. Colors indicate the different sites at which $C_{n}^{2}$ is observed with urban sites marked by (*).
Appendix C Details about baseline models